OpenStudy (anonymous):

The minimum value of (sinx+cosecx)^2 + (cosx+secx)^2

OpenStudy (anonymous):

Sin^2 x + 2 + cosec^2 x + cos^2 x + 2 + sec^2 x 5 + cosec^2 x + sec^2 x

OpenStudy (anonymous):

\[\sin^2x+2+\frac{ 1 }{ \sin^2x }+\cos^2x+2+\frac{ 1 }{ \cos^2x }=1+4+\frac{ 1 }{ (\sin^2x)(\cos^2x) }\] \[1+4+1=6\]

OpenStudy (anonymous):

minimum 6

OpenStudy (anonymous):

nope...the answer shuld be 9

OpenStudy (shubhamsrg):

yep its 9 only,, @Jonask just missed this: 1/sin^2x cos^2x = 4/(sin2x)^2 => minima is 4 hence total = 1 + 4 +4 = 9

OpenStudy (anonymous):

Hw Did u Do that @shubhamsrg

OpenStudy (anonymous):

\[\frac{ 1 }{ (\frac{ 1 }{ 2 }\sin2x)^2 }=\frac{ 4 }{ \sin^22x }\]

OpenStudy (anonymous):

min \[\frac{ 4 }{ 1 }=4\] instead of 1+4+1 1+4+4

OpenStudy (shubhamsrg):

anyways, you may also go for AM>= GM (sinx+cosecx)^2 + (cosx+secx)^2 >= 2 (sinx+cosecx) (cosx+secx) consider RHS 2 (sinx cosx + sinx sec x + cosecx cosx + cosecx sec x) 2( sinx cos x+ 1/sinx cosx + tanx + cot x) sin 2x + 4/sin 2x + 2tanx + 2/tanx x=pi/4 gives a maximum = 1 + 4 + 2+ 2 = 9 which is a minima for LHS..